# Calogero–Sutherland hyperbolic system and Heckman–Opdam $\mathfrak{gl}_{n}$ hypergeometric function

### v1

@@ A.3 Proofs of lemmas
$\begin{split}e^{x}-1\geq x,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ x<0,\end{split}$ (A.97)
@@
$\begin{split}\frac{\partial^{2}\beta_{g}(x)}{\partial x\partial g}\geq\frac{\partial^{2}\beta_{g}(x)}{\partial x\partial g}\bigg{|}_{g=1}\geq 0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ g\geq 1,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ x<0.\end{split}$ (A.101)
@@
$(g-1)(1-e^{t})U(t)\geq 0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ t<0,$ (A.105)
@@
$(g-1)(-t)(1-e^{t})U(t)\leq(g-1)^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ t<0.$ (A.107)
@@
According to (A.105) and (A.107), we know that $\sup_{t\in(-\infty,-1)}|tU(t)|\leq B$ for $B$ not depending on the choice of $\omega_{2}$. Also, due to (A.106) we know asymptotics of $U(t)$ at -$\infty$:
@@
$\begin{split}&\int^{-1}_{-\pi y^{-1}/2}\cos(yt)U(t)dt\leq\int^{-1}_{-\pi y^{-1}/2}\frac{(g-1)dt}{(-t)(1-e^{t})}=(g-1)\int^{-1}_{-\pi y^{-1}/2}\frac{dt}{-t}\\ &+(g-1)\int^{-1}_{-\pi y^{-1}/2}\frac{dt}{(-t)(e^{-t}-1)}\leq(g-1)(-\log(|t|))\Big{|}^{-1}_{-\pi y^{-1}/2}\\ &+(g-1)\int^{-1}_{-\infty}\frac{dt}{(-t)(e^{-t}-1)}=(g-1)\log(\pi y^{-1}/2)+D=(1-g)\log y+\widehat{D}\end{split}$ (A.114)
@@ Appendix B Asymptotics of wave function
$\begin{split}&\hat{\Psi}_{\bm{\lambda}_{n}}(\bm{x}_{n})=e^{2\pi\imath x_{n}(\lambda_{1}+\ldots+\lambda_{n})}\sum^{n-1}_{r=0}\hat{d}_{n-1-r}\int_{(\mathbb{R}+\imath\frac{g}{2}+\imath\varepsilon)^{n-1-r}}d\bm{t}_{n-1-r}\\[6.0pt] &\times\sum_{\begin{subarray}{c}J\subset{\{1,\ldots,n\}}\\ |J|=r\end{subarray}}\hat{\Psi}_{(t_{1},\ldots,t_{n-1-r};\{\lambda_{i}+\imath g/2\}_{i\in J})}(\bm{x}_{n-1}-x_{n}\bm{1})\prod_{\begin{subarray}{c}j\notin J\\ k\in J\end{subarray}}\hat{K}(\lambda_{j}-\lambda_{k}-\imath g/2)\\[3.0pt] &\times\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}(\lambda_{j}-t_{l})\prod_{\begin{subarray}{c}j\neq k\\ 1\leq j,k\leq n-1-r\end{subarray}}\hat{w}(t_{j}-t_{k})\prod_{\begin{subarray}{c}j\in J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}^{-1}(\lambda_{j}-t_{l}+\imath g).\end{split}$ (B.1)
@@
$\begin{split}\bigl{|}\check{\Phi}_{\bm{\lambda}_{n}}(\bm{x}_{n})\bigr{|}\leq C(g)\,\prod_{i<j}(1+|\mathrm{Re}\,\lambda_{i}-\mathrm{Re}\,\lambda_{j}|)^{N(n)}\,e^{-2\pi(x_{1}\operatorname{Im}\lambda_{1}+\ldots+x_{n}\operatorname{Im}\lambda_{n})},\end{split}$ (B.3)
@@
$C(g,\varepsilon)\,e^{\pi g\sum_{1\leq i\leq n-1}|x_{i}-x_{n}|}\,\prod_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}(1+|\mathrm{Re}\,(\lambda_{j}-\lambda_{k})|)^{1-g}\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}(1+|\mathrm{Re}\,\lambda_{j}-t_{l}|)^{g-1}\\ \times\hat{\Delta}(\bm{t}_{n-1-r})^{1-g}\,\exp\pi\bigg{[}\sum_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}|\mathrm{Re}\,(\lambda_{j}-\lambda_{k})|+\sum_{1\leq j<k\leq n-1-r}|t_{j}-t_{k}|-\sum_{j\notin J}\sum_{l=1}^{n-1-r}|\mathrm{Re}\,\lambda_{j}-t_{l}|\bigg{]}\\ \times\big{|}\check{\Phi}_{(\bm{t}_{n-1-r}+(\imath g/2+\imath\varepsilon)\bm{1};\leavevmode\nobreak\ \{\lambda_{i}+\imath g/2\}_{i\in J})}(\bm{x}_{n-1}-x_{n}\bm{1})\big{|},$ (B.5)
@@
Observe that Corollaries B.1B.2 are exactly Theorems 5.25.1 from Section 5.1.
@@ B.1 Proofs of lemmas
$\begin{split}&I_{r,J}(\bm{x}_{n};{\bm{\lambda}}_{n};\bm{t}_{n-1};\varepsilon)\\[6.0pt] &=e^{2\pi\imath x_{n}\underline{\bm{\lambda}}_{n}}\hat{d}_{n-1-r}\int_{\mathbb{R}^{n-1-r}}d\bm{t}_{n-1-r}\;\hat{\Psi}_{(\bm{t}_{n-1-r}+(\imath g/2+\imath\varepsilon)\,\bm{1};\{\lambda_{i}+\imath g/2\}_{i\in J})}(\bm{x}_{n-1}-x_{n}\bm{1})\\[6.0pt] &\times\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}(\lambda_{j}-t_{l}-\imath g/2-\imath\varepsilon)\prod_{\begin{subarray}{c}j\neq k\\ 1\leq j,k\leq n-1-r\end{subarray}}\hat{w}(t_{j}-t_{k})\\[6.0pt] &\times\prod_{\begin{subarray}{c}j\in J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}^{-1}(\lambda_{j}-t_{l}+\imath g/2-\imath\varepsilon)\,\prod_{\begin{subarray}{c}j\notin J\\ k\in J\end{subarray}}\hat{K}(\lambda_{j}-\lambda_{k}-\imath g/2).\end{split}$ (B.14)
@@
$\begin{split}F_{1}(\bm{x}_{n};\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)&=e^{\pi g\sum_{1\leq i<j\leq n}|x_{i}-x_{j}|}\prod_{1\leq i<j\leq n}\hat{w}(\lambda_{i}-\lambda_{j})\\[6.0pt] &\times e^{-\pi g\sum_{1\leq i<j\leq n-1}|x_{i}-x_{j}|}\,\prod_{1\leq i<j\leq n-1-r}\hat{w}(t_{i}-t_{j})^{-1}\\[6.0pt] &\times\prod_{\begin{subarray}{c}i<j\\ i,j\in J\end{subarray}}\hat{w}(\lambda_{i}-\lambda_{j})^{-1}\,\prod_{\begin{subarray}{c}i=1,\ldots,n-1-r\\ j\in J\end{subarray}}\hat{w}(t_{i}-\lambda_{j}+\imath\varepsilon)^{-1},\end{split}$ (B.15)
@@
$\begin{split}&F_{2}(\bm{x}_{n};\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)=\check{\Phi}_{(\bm{t}_{n-1-r}+(\imath g/2+\imath\varepsilon)\bm{1};\{\lambda_{i}+\imath g/2\}_{i\in J})}(\bm{x}_{n-1}-x_{n}\bm{1});\\[8.0pt] &F_{3}(\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)=\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}(\lambda_{j}-t_{l}-\imath g/2-\imath\varepsilon)\\[6.0pt] &\times\prod_{\begin{subarray}{c}j\neq k\\ 1\leq j,k\leq n-1-r\end{subarray}}\hat{w}(t_{j}-t_{k})\prod_{\begin{subarray}{c}j\in J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}^{-1}(\lambda_{j}-t_{l}+\imath g/2-\imath\varepsilon);\\[6.0pt] &F_{0}(x_{n};\bm{\lambda}_{n})=e^{2\pi\imath x_{n}\underline{\bm{\lambda}}_{n}}\hat{d}_{n-1-r};\qquad F_{4}(\bm{\lambda}_{n})=\prod_{\begin{subarray}{c}j\notin J\\ k\in J\end{subarray}}\hat{K}(\lambda_{j}-\lambda_{k}-\imath g/2).\end{split}$ (B.17)
@@
$\begin{split}\check{I}_{r,J}=F_{0}\int_{\mathbb{R}^{n-1-r}}d\bm{t}_{n-1-r}\;F_{2}(\bm{t}_{n-1-r})F_{3}(\bm{t}_{n-1-r})F_{4}({\bm{\lambda}}_{n})F_{1}(\bm{t}_{n-1-r}).\end{split}$ (B.18)
@@
$\begin{split}&\Big{|}F_{1}(\bm{x}_{n};\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)F_{3}(\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)F_{4}(\bm{\lambda}_{n})\Big{|}\leq C(g,\varepsilon)\,e^{\pi g\sum_{1\leq i\leq n-1}|x_{i}-x_{n}|}\\[6.0pt] &\times\prod_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}(1+|\mathrm{Re}\,(\lambda_{j}-\lambda_{k})|)^{1-g}\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}(1+|\mathrm{Re}\,\lambda_{j}-t_{l}|)^{g-1}\,\hat{\Delta}(\bm{t}_{n-1-r})^{1-g}\\[6.0pt] &\times\exp\pi\bigg{\{}\sum_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}|\mathrm{Re}\,(\lambda_{j}-\lambda_{k})|+\sum_{1\leq j<k\leq n-1-r}|t_{j}-t_{k}|-\sum_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}|\mathrm{Re}\,\lambda_{j}-t_{l}|\bigg{\}}.\end{split}$ (B.20)
@@
$\begin{split}e^{-2\pi\varepsilon\sum_{1\leq i\leq n-1-r}(x_{r+i}-x_{n})}\exp\bigg{\{}-2\pi\sum_{j_{1}<\ldots<j_{r}}\sum_{1\leq i\leq r}(x_{i}-x_{n})\operatorname{Im}\lambda_{j_{i}}\bigg{\}}\\[6.0pt] \leq e^{-2\pi\sum_{1\leq i\leq n-1}(x_{i}-x_{n})\operatorname{Im}\lambda_{i}},\end{split}$ (B.22)
@@
$\begin{split}\big{|}\check{I}_{r,J}\big{|}\leq C_{r,J}(g)\,\hat{\Delta}(\mathrm{Re}\,{\bm{\lambda}}_{n})^{N(n)}\,e^{-2\pi(x_{1}\operatorname{Im}\lambda_{1}+\ldots+x_{n}\operatorname{Im}\lambda_{n})}.\end{split}$ (B.23)
Consider the case r=n1r=n-1. By the formula (LABEL:int_pres) with $J=\{1,\ldots,n-1\}$
$\begin{split}\big{|}I_{n-1,\{1,\ldots,n-1\}}\big{|}=\Big{|}e^{2\pi\imath x_{n}(\lambda_{1}+\ldots+\lambda_{n})}\hat{\Psi}_{(\lambda_{1}+\imath g/2,\ldots,\lambda_{n-1}+\imath g/2)}(\bm{x}_{n-1}-x_{n}\bm{1})\\[6.0pt] \times\prod_{1\leq i\leq n-1}\hat{K}(\lambda_{n}-\lambda_{i}-\imath g/2)\Big{|}\leq e^{-2\pi x_{n}\sum_{i}\operatorname{Im}\lambda_{i}}e^{-\pi g\sum_{1\leq i\leq n-1}(x_{i}-x_{n})}\\[6.0pt] \times\big{|}\hat{\Psi}_{(\lambda_{1},\ldots,\lambda_{n-1})}(\bm{x}_{n-1}-x_{n}\bm{1})\big{|}\,\prod_{1\leq i\leq n-1}\big{|}\hat{K}(\lambda_{n}-\lambda_{i}-\imath g/2)\big{|}.\end{split}$ (B.24)
@@
$\begin{split}\frac{e^{\pi g\sum_{1\leq i<j\leq n}|x_{i}-x_{j}|}\prod_{1\leq i<j\leq n}\hat{w}(\lambda_{i}-\lambda_{j})}{e^{\pi g\sum_{1\leq i<j\leq n-1}|x_{i}-x_{j}|}\prod_{1\leq i<j\leq n-1}\hat{w}(\lambda_{i}-\lambda_{j})}.\end{split}$ (B.25)
@@
$\begin{split}\big{|}\check{I}_{n-1,\{1,\ldots,n-1\}}\big{|}\leq C_{r,J}(g)\,\hat{\Delta}(\mathrm{Re}\,{\bm{\lambda}}_{n-1})^{N(n-1)}\,e^{-2\pi\sum_{1\leq i\leq n}x_{i}\operatorname{Im}\lambda_{i}}.\end{split}$ (B.26)
@@
$\begin{split}|\check{\Phi}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})|\leq\sum_{J\subset\{1,\ldots,n\}}C_{|J|,J}(g)\,\hat{\Delta}(\mathrm{Re}\,{\bm{\lambda}}_{n})^{N(n)}\,e^{-2\pi\sum_{1\leq i\leq n}x_{i}\operatorname{Im}\lambda_{i}},\end{split}$ (B.27)
@@
$\begin{split}A=N(n-1),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ B=N(n-1),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C=N(n-1)+1-g\end{split}$ (B.28)
@@
$\begin{split}N(n)=A+2B+C+1=4N(n-1)+2-g.\end{split}$ (B.29)
@@
$\begin{split}N(n)=4^{n-2}g-\frac{(4^{n-2}-1)(g-2)}{3},\end{split}$ (B.30)
@@
$\begin{split}\big{|}F_{1}(\bm{x}_{n};\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)F_{3}(\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)F_{4}(\bm{\lambda}_{n})\big{|}\leq C(g,\varepsilon)\,e^{\pi g\sum_{1\leq i\leq n-1}|x_{i}-x_{n}|}\,\\[6.0pt] \times\prod_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}(1+|\lambda_{j}-\lambda_{k}|)^{1-g}\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}(1+|\lambda_{j}-t_{l}|)^{g-1}\,\hat{\Delta}(\bm{t}_{n-1-r})^{1-g}\\ \times\exp\pi\bigg{\{}\sum_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}|\lambda_{j}-\lambda_{k}|+\sum_{1\leq j<k\leq n-1-r}|t_{j}-t_{k}|-\sum_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}|\lambda_{j}-t_{l}|\bigg{\}}.\end{split}$ (B.31)

### v3

@@ A.3 Proofs of lemmas
$\begin{split}e^{x}-1\geq x,\penalty 10000\ \penalty 10000\ \penalty 10000\ x<0,\end{split}$ (A.49)
@@
$\begin{split}\frac{\partial^{2}\beta_{g}(x)}{\partial x\partial g}\geq\frac{\partial^{2}\beta_{g}(x)}{\partial x\partial g}\bigg|_{g=1}\geq 0,\penalty 10000\ \penalty 10000\ \penalty 10000\ g\geq 1,\penalty 10000\ \penalty 10000\ \penalty 10000\ x<0.\end{split}$ (A.53)
@@
$(g-1)(1-e^{t})U(t)\geq 0,\penalty 10000\ \penalty 10000\ \penalty 10000\ t<0,$ (A.57)
@@
$(g-1)(-t)(1-e^{t})U(t)\leq(g-1)^{2},\penalty 10000\ \penalty 10000\ \penalty 10000\ t<0.$ (A.59)
@@
According to (A.57) and (A.59), we know that
$\sup_{t\in(-\infty,-1)}|tU(t)|\leq B$
for $B$ not depending on the choice of $\omega_{2}$. Also, due to (A.58) we know asymptotics of $U(t)$ at -$\infty$:
@@
$\begin{split}&\int^{-1}_{-\pi y^{-1}/2}\cos(yt)U(t)dt\leq\int^{-1}_{-\pi y^{-1}/2}\frac{(g-1)dt}{(-t)(1-e^{t})}=(g-1)\int^{-1}_{-\pi y^{-1}/2}\frac{dt}{-t}\\ &+(g-1)\int^{-1}_{-\pi y^{-1}/2}\frac{dt}{(-t)(e^{-t}-1)}\leq(g-1)(-\log(|t|))\Big|^{-1}_{-\pi y^{-1}/2}\\ &+(g-1)\int^{-1}_{-\infty}\frac{dt}{(-t)(e^{-t}-1)}\\ &=(g-1)\log(\pi y^{-1}/2)+D=(1-g)\log y+\widehat{D}\end{split}$ (A.66)
@@ Appendix B Asymptotics of wave function
$\begin{split}&\hat{\Psi}_{\bm{\lambda}_{n}}(\bm{x}_{n})=e^{2\pi\imath x_{n}(\lambda_{1}+\ldots+\lambda_{n})}\sum^{n-1}_{r=0}\hat{d}_{n-1-r}\int_{(\mathbb{R}+\imath\frac{g}{2}+\imath\varepsilon)^{n-1-r}}d\bm{t}_{n-1-r}\times\\[6.0pt] &\sum_{\begin{subarray}{c}J\subset{\{1,\ldots,n\}}\\ |J|=r\end{subarray}}\hat{\Psi}_{(t_{1},\ldots,t_{n-1-r};\{\lambda_{i}+\imath g/2\}_{i\in J})}(\bm{x}_{n-1}-x_{n}\bm{1})\prod_{\begin{subarray}{c}j\notin J\\ k\in J\end{subarray}}\hat{K}(\lambda_{j}-\lambda_{k}-\imath g/2)\times\\[3.0pt] &\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}\!\!\!\hat{K}(\lambda_{j}-t_{l})\!\!\prod_{\begin{subarray}{c}j\neq k\\ 1\leq j,k\leq n-1-r\end{subarray}}\hat{w}(t_{j}-t_{k})\prod_{\begin{subarray}{c}j\in J\\ l=1,\ldots,n-1-r\end{subarray}}\!\!\!\hat{K}^{-1}(\lambda_{j}-t_{l}+\imath g).\end{split}$ (B.1)
@@
$\begin{split}\bigl|\check{\Phi}_{\bm{\lambda}_{n}}(\bm{x}_{n})\bigr|\leq C(g)\,\prod_{i<j}(1+|\mathrm{Re}\,\lambda_{i}-\mathrm{Re}\,\lambda_{j}|)^{N(n)}\,e^{-2\pi(x_{1}\operatorname{Im}\lambda_{1}+\ldots+x_{n}\operatorname{Im}\lambda_{n})},\end{split}$
@@
$C(g,\varepsilon)\,e^{\pi g\!\!\!\!\!\sum\limits_{1\leq i\leq n-1}\!\!|x_{i}-x_{n}|}\,\!\!\prod_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}(1+|\mathrm{Re}\,(\lambda_{j}-\lambda_{k})|)^{1-g}\!\!\!\!\!\!\!\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}\!\!\!\!\!\!\!(1+|\mathrm{Re}\,\lambda_{j}-t_{l}|)^{g-1}\\ \times\,\exp\pi\bigg[\sum_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}|\mathrm{Re}\,(\lambda_{j}-\lambda_{k})|+\sum_{1\leq j<k\leq n-1-r}|t_{j}-t_{k}|-\sum_{j\notin J}\sum_{l=1}^{n-1-r}|\mathrm{Re}\,\lambda_{j}-t_{l}|\bigg]\\ \times\hat{\Delta}(\bm{t}_{n-1-r})^{1-g}\big|\check{\Phi}_{(\bm{t}_{n-1-r}+(\imath g/2+\imath\varepsilon)\bm{1};\penalty 10000\ \{\lambda_{i}+\imath g/2\}_{i\in J})}(\bm{x}_{n-1}-x_{n}\bm{1})\big|,$ (B.4)
@@
Observe that Corollaries B.1, B.2 are exactly Theorems 5.2, 5.1 from Section 5.1.
@@ B.1 Proofs of lemmas
$\begin{split}&I_{r,J}(\bm{x}_{n};{\bm{\lambda}}_{n};\bm{t}_{n-1};\varepsilon)=e^{2\pi\imath x_{n}\underline{\bm{\lambda}}_{n}}\hat{d}_{n-1-r}\\[6.0pt] &\times\int_{\mathbb{R}^{n-1-r}}d\bm{t}_{n-1-r}\;\hat{\Psi}_{(\bm{t}_{n-1-r}+(\imath g/2+\imath\varepsilon)\,\bm{1};\{\lambda_{i}+\imath g/2\}_{i\in J})}(\bm{x}_{n-1}-x_{n}\bm{1})\\[6.0pt] &\times\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}(\lambda_{j}-t_{l}-\imath g/2-\imath\varepsilon)\prod_{\begin{subarray}{c}j\neq k\\ 1\leq j,k\leq n-1-r\end{subarray}}\hat{w}(t_{j}-t_{k})\\[6.0pt] &\times\prod_{\begin{subarray}{c}j\in J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}^{-1}(\lambda_{j}-t_{l}+\imath g/2-\imath\varepsilon)\,\prod_{\begin{subarray}{c}j\notin J\\ k\in J\end{subarray}}\hat{K}(\lambda_{j}-\lambda_{k}-\imath g/2).\end{split}$ (B.13)
@@
$\begin{split}F_{1}(\bm{x}_{n};&\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)=e^{\pi g\sum_{1\leq i<j\leq n}|x_{i}-x_{j}|}\prod_{1\leq i<j\leq n}\hat{w}(\lambda_{i}-\lambda_{j})\\[6.0pt] &\times e^{-\pi g\sum_{1\leq i<j\leq n-1}|x_{i}-x_{j}|}\,\prod_{1\leq i<j\leq n-1-r}\hat{w}(t_{i}-t_{j})^{-1}\\[6.0pt] &\times\prod_{\begin{subarray}{c}i<j\\ i,j\in J\end{subarray}}\hat{w}(\lambda_{i}-\lambda_{j})^{-1}\,\prod_{\begin{subarray}{c}i=1,\ldots,n-1-r\\ j\in J\end{subarray}}\hat{w}(t_{i}-\lambda_{j}+\imath\varepsilon)^{-1},\end{split}$ (B.14)
@@
$\begin{split}&F_{2}(\bm{x}_{n};\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)=\check{\Phi}_{(\bm{t}_{n-1-r}+(\imath g/2+\imath\varepsilon)\bm{1};\{\lambda_{i}+\imath g/2\}_{i\in J})}(\bm{x}_{n-1}-x_{n}\bm{1});\\[8.0pt] &F_{3}(\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)=\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}(\lambda_{j}-t_{l}-\imath g/2-\imath\varepsilon)\\[6.0pt] &\times\prod_{\begin{subarray}{c}j\neq k\\ 1\leq j,k\leq n-1-r\end{subarray}}\hat{w}(t_{j}-t_{k})\prod_{\begin{subarray}{c}j\in J\\ l=1,\ldots,n-1-r\end{subarray}}\hat{K}^{-1}(\lambda_{j}-t_{l}+\imath g/2-\imath\varepsilon);\\[6.0pt] &F_{0}(x_{n};\bm{\lambda}_{n})=e^{2\pi\imath x_{n}\underline{\bm{\lambda}}_{n}}\hat{d}_{n-1-r};\qquad F_{4}(\bm{\lambda}_{n})=\prod_{\begin{subarray}{c}j\notin J\\ k\in J\end{subarray}}\hat{K}(\lambda_{j}-\lambda_{k}-\imath g/2).\end{split}$
@@
$\begin{split}\check{I}_{r,J}=F_{0}\int_{\mathbb{R}^{n-1-r}}d\bm{t}_{n-1-r}\;F_{2}(\bm{t}_{n-1-r})F_{3}(\bm{t}_{n-1-r})F_{4}({\bm{\lambda}}_{n})F_{1}(\bm{t}_{n-1-r}).\end{split}$
@@
$\begin{split}&\Big|F_{1}(\bm{x}_{n};\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)F_{3}(\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)F_{4}(\bm{\lambda}_{n})\Big|\leq C(g,\varepsilon)\,e^{\pi g\sum\limits_{1\leq i\leq n-1}|x_{i}-x_{n}|}\\[6.0pt] &\times\prod_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}(1+|\mathrm{Re}\,(\lambda_{j}-\lambda_{k})|)^{1-g}\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}(1+|\mathrm{Re}\,\lambda_{j}-t_{l}|)^{g-1}\,\hat{\Delta}(\bm{t}_{n-1-r})^{1-g}\times\\[6.0pt] &\exp\pi\bigg\{\sum_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}|\mathrm{Re}\,(\lambda_{j}-\lambda_{k})|+\sum_{1\leq j<k\leq n-1-r}|t_{j}-t_{k}|-\sum_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}|\mathrm{Re}\,\lambda_{j}-t_{l}|\bigg\}.\end{split}$ (B.17)
@@
$\begin{split}e^{-2\pi\varepsilon\sum_{1\leq i\leq n-1-r}(x_{r+i}-x_{n})}\exp\bigg\{-2\pi\sum_{j_{1}<\ldots<j_{r}}\sum_{1\leq i\leq r}(x_{i}-x_{n})\operatorname{Im}\lambda_{j_{i}}\bigg\}\\[6.0pt] \leq e^{-2\pi\sum_{1\leq i\leq n-1}(x_{i}-x_{n})\operatorname{Im}\lambda_{i}},\end{split}$
@@
$\begin{split}\big|\check{I}_{r,J}\big|\leq C_{r,J}(g)\,\hat{\Delta}(\mathrm{Re}\,{\bm{\lambda}}_{n})^{N(n)}\,e^{-2\pi(x_{1}\operatorname{Im}\lambda_{1}+\ldots+x_{n}\operatorname{Im}\lambda_{n})}.\end{split}$
Consider the case r=n1r=n-1. By (LABEL:int_pres) with $J=\{1,\ldots,n-1\}$
$\begin{split}\big|I_{n-1,\{1,\ldots,n-1\}}\big|=\Big|e^{2\pi\imath x_{n}(\lambda_{1}+\ldots+\lambda_{n})}\hat{\Psi}_{(\lambda_{1}+\imath g/2,\ldots,\lambda_{n-1}+\imath g/2)}(\bm{x}_{n-1}-x_{n}\bm{1})\\[6.0pt] \times\prod_{1\leq i\leq n-1}\hat{K}(\lambda_{n}-\lambda_{i}-\imath g/2)\Big|\leq e^{-2\pi x_{n}\sum_{i}\operatorname{Im}\lambda_{i}}e^{-\pi g\sum_{1\leq i\leq n-1}(x_{i}-x_{n})}\\[6.0pt] \times\big|\hat{\Psi}_{(\lambda_{1},\ldots,\lambda_{n-1})}(\bm{x}_{n-1}-x_{n}\bm{1})\big|\,\prod_{1\leq i\leq n-1}\big|\hat{K}(\lambda_{n}-\lambda_{i}-\imath g/2)\big|.\end{split}$
@@
$\begin{split}\frac{e^{\pi g\sum_{1\leq i<j\leq n}|x_{i}-x_{j}|}\prod_{1\leq i<j\leq n}\hat{w}(\lambda_{i}-\lambda_{j})}{e^{\pi g\sum_{1\leq i<j\leq n-1}|x_{i}-x_{j}|}\prod_{1\leq i<j\leq n-1}\hat{w}(\lambda_{i}-\lambda_{j})}.\end{split}$
@@
$\begin{split}\big|\check{I}_{n-1,\{1,\ldots,n-1\}}\big|\leq C_{r,J}(g)\,\hat{\Delta}(\mathrm{Re}\,{\bm{\lambda}}_{n-1})^{N(n-1)}\,e^{-2\pi\sum_{1\leq i\leq n}x_{i}\operatorname{Im}\lambda_{i}}.\end{split}$
@@
$\begin{split}|\check{\Phi}_{{\bm{\lambda}}_{n}}(\bm{x}_{n})|\leq\sum_{J\subset\{1,\ldots,n\}}C_{|J|,J}(g)\,\hat{\Delta}(\mathrm{Re}\,{\bm{\lambda}}_{n})^{N(n)}\,e^{-2\pi\sum_{1\leq i\leq n}x_{i}\operatorname{Im}\lambda_{i}},\end{split}$
@@
$\begin{split}A=N(n-1),\penalty 10000\ \penalty 10000\ \penalty 10000\ B=N(n-1),\penalty 10000\ \penalty 10000\ \penalty 10000\ C=N(n-1)+1-g\end{split}$
@@
$\begin{split}N(n)=A+2B+C+1=4N(n-1)+2-g.\end{split}$
@@
$\begin{split}N(n)=4^{n-2}g-\frac{(4^{n-2}-1)(g-2)}{3},\end{split}$
@@
$\begin{split}\big|F_{1}(\bm{x}_{n};\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)F_{3}(\bm{\lambda}_{n};\bm{t}_{n-1-r};\varepsilon)F_{4}(\bm{\lambda}_{n})\big|\leq C(g,\varepsilon)\,e^{\pi g\sum\limits_{1\leq i\leq n-1}|x_{i}-x_{n}|}\,\\[6.0pt] \times\prod_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}(1+|\lambda_{j}-\lambda_{k}|)^{1-g}\prod_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}(1+|\lambda_{j}-t_{l}|)^{g-1}\,\hat{\Delta}(\bm{t}_{n-1-r})^{1-g}\\ \times\exp\pi\bigg\{\sum_{\begin{subarray}{c}j<k\\ j,k\notin J\end{subarray}}|\lambda_{j}-\lambda_{k}|+\sum_{1\leq j<k\leq n-1-r}|t_{j}-t_{k}|-\sum_{\begin{subarray}{c}j\notin J\\ l=1,\ldots,n-1-r\end{subarray}}|\lambda_{j}-t_{l}|\bigg\}.\end{split}$ (B.19)